# Anomalous dimensions from conformal field theory: generalized $\phi^{2n+1}$ theories

### v1

@@ 1 Introduction
After discussing the $n=1$ case, we proceed to consider theories with higher $n$ [^9] For studies of the canonical $n=2$ theory with $S_{N+1}$ symmetry, see Codello:2018nbe ; Codello:2020mnt . We only consider the single-field case for the $n>1$ theory in this paper. For $n=2$, the interaction is associated with $\phi^{5}$, and the corresponding theory is related to the Blume-Capel model. More broadly, we study the generalized $\phi^{2n+1}$ theory with an integer $n>1$. A notable feature is that, at leading order in the $\epsilon$ expansion, a family of data is consistent with conformal symmetry, as far as our analysis is concerned. This means that a large set of leading anomalous dimensions is determined up to one parameter, in contrast to the $n=1$ case where the anomalous dimensions are fixed. One motivation for considering the $n>1$ cases is that these theories are associated with multicritical Yang-Lee models vonGehlen:1994rp ; Lencses:2022ira ; Lencses:2023evr ; Lencses:2024wib , and they provide a series of universality classes containing $\mathbb{Z}_{2}$-odd multicritical points. See also Zambelli:2016cbw ; Gracey:2017okb ; Codello:2017qek ; Codello:2017epp for related studies. This concludes our discussion of the theories that will be investigated in this paper.
@@
where the indices are suppressed for simplicity. The partially conserved currents do not vanish when we act with fewer derivatives. These operators are related to higher spin symmetry of the free theory. One can construct conserved charges from the partially conserved currents, generalizing the $k=1$ higher spin algebra. We refer to Brust:2016gjy for more details. For $k>1$, the partially conserved currents play a significant role in the analytic bootstrap, as their twists are lower than those of the usual conserved currents. In the lightcone limit, the contribution from the partially conserved currents or their interacting counterparts are more important than the Regge trajectory associated with the conserved currents.
To derive the anomalous dimensions of the bilinear operators, we employ the conformal multiplet recombination method initiated in Rychkov:2015naa , which considered the canonical $\phi^{4}$ theory. In Rychkov:2015naa , it was further suggested that the results can be extended to the canonical $\phi^{3}$ theory at $d=6-\epsilon$. A key difference is the absence of $\mathbb{Z}_{2}$ symmetry in the $\phi^{3}$ theory, in contrast to the $\phi^{4}$ case. Since the procedure in Rychkov:2015naa relies on the $\mathbb{Z}_{2}$ symmetry, its generalization to the $\mathbb{Z}_{2}$-odd cases does not seem straightforward. A similar approach based on the equation of motion was proposed to study the $\phi^{3}$ theory Nii:2016lpa . By applying the equation of motion to correlation functions, the constraints on the anomalous dimensions can be derived more readily. In this paper, we revisit the $\phi^{2n+1}$ theory from the multiplet recombination perspective and extend the results to some spinning operators. As a consistency check, we also verify that the results for the anomalous dimensions and OPE coefficients are consistent with crossing symmetry using the analytic bootstrap.
The multiplet recombination asserts that free multiplets recombine into long multiplets at interacting IR fixed points. For instance, at $d=4-\epsilon$, the Wilson-Fisher multiplet $\{\phi\}$ is a long multiplet into which the free multiplets $\{\phi_{\text{f}}\}$ and $\{\phi^{3}_{\text{f}}\}$ recombine. Here the subscript f implies the free theory. In the Gaussian limit, the descendant $\Box\phi$ is expected to become null and decouple, but a singular change in the normalization can lead to a physical operator, which is identified with the free theory primary operator $\phi^{3}_{\text{f}}$. In other words, the Wilson-Fisher counterpart of $\phi^{3}_{\text{f}}$ is identified as a descendant $\Box\phi$ up to a normalization factor  [^10] All physical operators in the free theory should have counterparts in the interacting theory.:
@@
Despite the resemblance to the equation of motion $\Box\phi\propto\epsilon\,\phi^{3}$, the multiplet recombination is conceptually different, as it can be formulated purely in the CFT language and does not require a Lagrangian description. The multiplet recombination above generalizes to various cases, including the $\mathbb{Z}_{2}$-odd deformations of the free higher-derivative CFTs Gliozzi:2016ysv ; Gliozzi:2017hni ; Gliozzi:2017gzh . See Basu:2015gpa ; Yamaguchi:2016pbj ; Hasegawa:2016piv ; Ghosh:2015opa ; Raju:2015fza ; Gliozzi:2016ysv ; Roumpedakis:2016qcg ; Soderberg:2017oaa ; Behan:2017emf ; Gliozzi:2017hni ; Gliozzi:2017gzh ; Nishioka:2022odm ; Nishioka:2022qmj ; Antunes:2022vtb ; Guo:2023qtt for applications of the multiplet recombination to numerous CFTs, and see Safari:2017irw ; Nii:2016lpa ; Hasegawa:2018yqg ; Skvortsov:2015pea ; Giombi:2016hkj ; Giombi:2017rhm ; Giombi:2016zwa ; Codello:2017qek ; Codello:2018nbe ; Antipin:2019vdg ; Vacca:2019rsh ; Giombi:2020rmc ; Giombi:2020xah ; Dey:2020jlc ; Giombi:2021cnr ; Safari:2021ocb ; Zhou:2022pah ; Bissi:2022bgu ; Herzog:2022jlx ; Giombi:2022vnz ; SoderbergRousu:2023pbe for applications with more emphasis on the equations of motion.
This paper is organized as follows. In section 2, we consider the generalized Yang-Lee edge singularity. We first employ the multiplet recombination to compute the anomalous dimension of $\phi$ and bilinear operators $\mathcal{J}$, then we use the analytic bootstrap to verify that the results are consist with crossing symmetry. In section 3, we carry out a similar analysis, obtaining the anomalous dimensions in the generalized Potts model, and we check that the results are consistent with crossing symmetry as well. Section 4 focuses on the generalized $\phi^{2n+1}$ theory with $n>1$, where we only consider the single-field case. We determine the anomalous dimensions up to a free parameter, then we verify the consistency between the multiplet-recombination results and crossing symmetry. We summarize the main results in table 1. In appendix A, we work out the action of Laplacians on the scalar three-point function. In appendix B, we verify, in all three cases, that the multiplet-recombination results are consistent with crossing symmetry at subleading twist. We briefly review some representation theory of the symmetric group in appendix C. More details about the tensor structures in the Potts model can be found in appendix D. We compute in appendix E the ratios of OPE coefficients relevant to the multiplet recombination. In appendix F, we derive an OPE coefficient used in the analytic bootstrap. Appendix G discusses some special limits in the generalized O($N$) symmetric $\phi^{4}$ theory.
@@
Table 1: Summary of the main results. The second column lists the anomalous dimensions of $\phi$, and the third column lists the anomalous dimensions of the bilinear operators of the schematic form $J\sim\phi\,\partial^{\ell}\Box^{m}\phi$.
@@ 2 Generalized Yang-Lee edge singularity
where the upper critical dimension is $d_{\text{u}}=6k$. The parameter $\mu$ has mass dimension 1, so the coupling constant $g$ is dimensionless. We emphasize that $g$ is allowed to be imaginary.
@@ 2.1 Multiplet recombination
$\langle\phi(x_{1})\phi(x_{2})\rangle=1/x_{12}^{\smash{2\Delta_{\phi}}}\,.$ (9)
@@ 2.2 Analytic bootstrap
In this subsection, we verify the consistency between the results above and crossing symmetry. In the lightcone limit of the crossing equation for the four-point function $\langle\phi\phi\phi\phi\rangle$, the anomalous dimensions of the bilinear operators are read off from the terms with enhanced singularity in the crossed channel. [^16] The meaning of the enhanced singularity is as follows. An expression with an enhanced singularity is power-law divergent after repeated action of the conformal Casimir operator Alday:2015eya . The results are consistent with those derived from the multiplet recombination.
@@ 3 Generalized Potts model
We can extend the $\phi^{3}$ discussion to the case of an $N$-component field $\phi_{a}$, where $a=1,2,\ldots,N$. The action that describes the ($N+1$)-state Potts model is
@@ 3.1 Multiplet recombination
This expression is very similar to (22), but here an $N$-dependent factor appears in the second term in the denominator. [^19] Interestingly, the anomalous dimension of $\phi$ is $k$-independent to order $\epsilon$ at $N=2$: $\gamma_{\phi}=\epsilon/6+O(\epsilon^{3/2})$. At $k=1$, the result agrees with that from the RG analysis Amit:1976pz . We also obtain $\alpha$ and the OPE coefficient of $\phi$:
@@
On the highest trajectory $m=k-1$, the spin-2 current in the trivial representation $(N+1)$ is the stress tensor. As expected, the anomalous dimension of the stress tensor is $O(\epsilon^{3/2})$ according to (69). At $k=1$, the set of leading anomalous dimensions $\{\gamma_{\phi},\gamma_{\mathcal{J}}\}$ contains negative elements when $N>7/3$, which indicates that the canonical Potts model is non-unitary near the upper critical dimension. Moreover, there exists a shadow relation in the standard representation:
$\Delta_{\smash{\phi}}+\Delta_{\mathcal{J}_{\text{standard rep}}}\big{|}_{\ell=m=0}=d+O(\epsilon^{3/2})\,,$ (71)
which is true for any positive integer $k$. In the $N\rightarrow\infty$ limit, this equation is related to (30) in the generalized Yang-Lee theory.
Below, we examine the Chew-Frautschi plots to gain further insights into the theory. The analytic continuation of the trajectories is given by Caron-Huot:2022eqs
@@
Figure 1: Chew-Frautschi plot for $\mathcal{J}^{(m=0)}_{\ell,\,\text{trivial rep}}$ at $N=3$. The dashed lines indicate the the $\epsilon\rightarrow 0$ limit of the Regge trajectory. We have set $k=1$ and $\epsilon=10^{-3}$.
Figure 2: Regge trajectories of $\mathcal{J}^{(m=0)}_{\ell\,\text{trivial rep}}$. The $N=\infty$ case corresponds to the decoupled Yang-Lee theory. The plot is made at $k=1$ and $\epsilon=10^{-3}$.
Figure 3: Regge trajectories associated with $\mathcal{J}^{(m=0)}_{\ell,\,\text{trivial rep}}$. The trajectory in the interacting theory approaches the free theory trajectory as $N\rightarrow 1$. We have used $k=1$ and $\epsilon=10^{-3}$.
@@
$\Delta^{*}_{\pm}-d/2=\pm\left(1+\frac{3R^{\prime}-11N+23}{6(3N-7)}\,\epsilon+O(\epsilon^{3/2})\right)\,,$ (73)
where we have defined $R^{\prime}\equiv R/(N+1)^{2}$. The solution $\Delta^{*}_{-}$ in the standard representation is precisely the scaling dimension of $\phi_{a}$, and $\Delta^{*}_{+}$ is given by $\Delta_{\mathcal{J_{\text{standard rep}}}}|_{\ell=m=0}$. This is consistent with the shadow relation (71), since $\Delta^{*}_{\pm}$ are related by the shadow transformation $\Delta\rightarrow d-\Delta$.
Figure 4: An $\mathbb{R}^{3}$ projection of the $\mathbb{C}^{2}$ Chew-Frautschi plot for $\mathcal{J}^{(m=0)}_{\ell,\,\text{trivial rep}}$ at $N=3$. The surfaces are the Riemann surfaces of the functions $\ell(\Delta)$. The imaginary part of $\ell$ is shown by color. Red regions correspond to $\text{Im}[\ell]>0$, and blue regions are associated with $\text{Im}[\ell]<0$. The two turquoise points represent two solutions to $\ell=0$, which are $\Delta_{\pm}$ in (73). The plot is symmetric with respect to the plane $\text{Re}[\Delta-d/2]=0$. We have chosen the parameters $k=1$ and $\epsilon=0.3$.
Figure 5: An $\mathbb{R}^{3}$ projection of the $\mathbb{C}^{2}$ Regge trajectories of $\mathcal{J}^{(m=0)}_{\ell,\,\text{trivial rep}}$ at $N=3$. The peaks at $\text{Re}[\ell]=-2$ and $\text{Im}[\ell]=0$ correspond to the $1/(\ell+2)$ pole in the anomalous dimension. We indicate the imaginary part of $\Delta-d/2$ by color. The regions with $\text{Im}[\Delta-d/2]>0$ are in red, and the regions with $\text{Im}[\Delta-d/2]<0$ are in blue. Two of the solutions to $\Delta=d/2$ are $\ell_{\pm}$ in (74). They form a complex conjugate pair, and are indicated by the two turquoise points. We have used $k=1$ and $\epsilon=0.3$.
Figure 6: A 3D Chew-Frautschi plot for $\mathcal{J}^{(m=0)}_{\ell,\,\text{trivial rep}}$ at $N=6/5$. Color coding indicates the imaginary part of $\ell$: red for $\text{Im}[\ell]>0$ and blue for $\text{Im}[\ell]<0$. The turquoise points are the solutions to $\ell=0$, given by $\Delta_{\pm}$ in (73). The plot is symmetric relative to the plane $\text{Re}[\Delta-d/2]=0$. The parameters are $k=1$ and $\epsilon=0.3$.
@@
Figure 7: An $\mathbb{R}^{3}$ projection of the $\mathbb{C}^{2}$ Chew-Frautschi plot for $\mathcal{J}^{(m=0)}_{\ell,\,\text{trivial rep}}$ at $N=6/5$. The peaks at $\text{Re}[\ell]=-2$, $\text{Im}[\ell]=0$ is due to the $1/(\ell+2)$ pole in the anomalous dimension. The imaginary part of $\ell$ is shown by color: red regions implies $\text{Im}[\ell]>0$, and blue regions indicate $\text{Im}[\ell]<0$. The turquoise points mark the solutions $\ell_{\pm}^{*}$ in (74). The plot is made at $k=1$ and $\epsilon=0.3$.
@@
Now, we briefly comment on the $k>1$ Chew-Frautschi plots. One difference is the presence of more curves, which arise from the deformation of more horizontal lines in the Gaussian limit. Another difference is that we can consider trajectories with higher $m$. Let us first discuss the Gaussian limit, where the scaling dimensions do not depend on the $S_{N+1}$ representations. For a positive integer $k$, the horizontal lines are given by $\frac{(1-3k-\ell-m)_{2k}}{\ell+k+2m}=0$, implying that there are $12k-$ 1 such lines. The diagonal lines are translations of those in the $k=1$ case. In the interacting theory, the trajectories are deformed into disconnected pieces. For example, the $\mathcal{J}^{(m=0)}_{\ell,\,\text{trivial rep}}$ trajectory in the $k=2$ theory and its Gaussian limit are shown in figure 8. As in the $k=1$ case, the Chew-Frautschi plots in the interacting theory depend on $N$. For both $m=0$ and $m=1$, there are three scenarios: $N>65/37$, $371<N<65/$ 37 and $N<1$. We present the Chew-Frautschi plots with different values of $N$ in figures 9, 10, 11 and 12. The Chew-Frautschi plots with $N<1$ are similar to case of $N>65/37$.
Figure 8: Chew-Frautschi plot for $\mathcal{J}^{(m=0)}_{\ell,\,\text{trivial rep}}$ in the $k=2$ theory. The interacting case with $N=2$ is shown in red. In the Gaussian limit, there are two dashed diagonal lines given by $(\ell+2)^{2}=(\Delta-6)^{2}$, and there exist three dashed horizontal lines at $\ell=-3,-4,-5$.
Figure 9: In the $k=2$ theory, the Regge trajectories of $\mathcal{J}^{(m=0)}_{\ell}$ are shown. These trajectories correspond to the case of $N>65/37$. We have set $\epsilon=10^{-3}$.
Figure 10: Chew-Frautschi plot of the $k=2$ interacting theory. The Regge trajectories of $\mathcal{J}^{(m=0)}_{\ell,\,\text{trivial rep}}$ with $371<N<65/$ 37 are shown. We have used $\epsilon=10^{-3}$.
Figure 11: Regge trajectories of $\mathcal{J}^{(m=1)}_{\ell,\,\text{trivial rep}}$ in the $k=2$ interacting theory. We present the scenario where $N>65/37$. Again, $\epsilon=10^{-3}$ is used.
Figure 12: In the $k=2$ interacting theory, we show the Regge trajectories of $\mathcal{J}^{(m=1)}_{\ell,\,\text{trivial rep}}$ with $371<N<65/$ 37, where $\epsilon$ is set to $10^{-3}$.

### v2

@@ 1 Introduction
After discussing the $n=1$ case, we proceed to consider theories with greater integer $n$ [^9] For studies of the canonical $n=2$ theory with $S_{N+1}$ symmetry, see Codello:2018nbe ; Codello:2020mnt . We only consider the single-field case for the $n>1$ theory in this paper. For $n=2$, the interaction is associated with $\phi^{5}$, and the corresponding CFT is expected to be a multicritical generalization of the Yang-Lee CFT. More broadly, we consider the generalized $\phi^{2n+1}$ theory with $n>1$. A notable feature is that, at leading order in the $\epsilon$ expansion, a family of data is consistent with conformal symmetry, as far as our analysis is concerned. This means that a large set of leading anomalous dimensions is determined up to one parameter, in contrast to the $n=1$ case where the anomalous dimensions are fixed. As the $\phi^{5}$ theory, the $\phi^{2n+1}$ theories with higher $n$ are also associated with multicritical Yang-Lee models vonGehlen:1994rp ; Lencses:2022ira ; Lencses:2023evr ; Lencses:2024wib , and they provide a series of universality classes containing multicritical points with $\mathbb{Z}_{2}$-odd interactions. See also Zambelli:2016cbw ; Gracey:2017okb ; Codello:2017qek ; Codello:2017epp for related studies. This concludes our discussion of the theories that will be investigated in this paper.
@@
where the indices are suppressed for simplicity. The partially conserved currents do not vanish when we act with fewer derivatives. These operators are related to higher spin symmetry of the free theory. One can construct conserved charges from the partially conserved currents, generalizing the $k=1$ higher spin algebra. We refer to Brust:2016gjy for more details. For $k>1$, the interacting versions of the partially conserved currents play a significant role in the analytic bootstrap, as their twists are lower than those of the usual conserved currents. In the lightcone limit, the interacting counterparts of the partially conserved currents are more important than the Regge trajectories associated with the conserved currents.
To derive the anomalous dimensions of the bilinear operators, we employ the conformal multiplet recombination method initiated in Rychkov:2015naa , which considered the canonical $\phi^{4}$ theory. In Rychkov:2015naa , it was further suggested that the results can be extended to the canonical $\phi^{3}$ theory at $d=6-\epsilon$.  [^10] According to Rychkov:2015naa ; Gliozzi:2016ysv ; Gliozzi:2017gzh , Yu Nakayama has considered the multiplet recombination for the $\epsilon=6-d$ expansion of the $\phi^{3}$ theory in his unpublished notes. A key difference is the absence of $\mathbb{Z}_{2}$ symmetry in the $\phi^{3}$ theory, in contrast to the $\phi^{4}$ case. Since the procedure in Rychkov:2015naa relies on the $\mathbb{Z}_{2}$ symmetry, its generalization to the $\mathbb{Z}_{2}$-odd cases does not seem straightforward. A similar approach based on the equation of motion was proposed to study the $\phi^{3}$ theory Nii:2016lpa . By applying the equation of motion to correlation functions, the constraints on the anomalous dimensions can be derived more readily. In this paper, we revisit the $\phi^{2n+1}$ theory from the multiplet recombination perspective and extend the results to some spinning operators. As a consistency check, we also verify that the results for the anomalous dimensions and OPE coefficients are consistent with crossing symmetry using the analytic bootstrap.
The multiplet recombination asserts that free multiplets recombine into long multiplets at interacting IR fixed points. For instance, at $d=4-\epsilon$, the Wilson-Fisher multiplet $\{\phi\}$ is a long multiplet, into which the free multiplets $\{\phi_{\text{f}}\}$ and $\{\phi^{3}_{\text{f}}\}$ recombine. Here the subscript f implies the free theory. In the Gaussian limit, the descendant $\Box\phi$ is expected to become null and decouple, but a singular change in the normalization can lead to a physical operator, which is identified with the free theory primary operator $\phi^{3}_{\text{f}}$. In other words, the Wilson-Fisher counterpart of $\phi^{3}_{\text{f}}$ is identified as a descendant $\Box\phi$ up to a normalization factor  [^11] All physical operators in the free theory should have counterparts in the interacting theory.:
@@
Despite the resemblance to the equation of motion $\Box\phi\propto\epsilon\,\phi^{3}$, the multiplet recombination is conceptually different, as it can be formulated purely in the CFT language and does not require a Lagrangian description. The multiplet recombination above generalizes to various cases. As in Guo:2023qtt , we follow a suggestion of Gliozzi:2017gzh and we apply the multiplet recombination method to the $\mathbb{Z}_{2}$-odd deformations of the free higher-derivative CFTs. See Basu:2015gpa ; Yamaguchi:2016pbj ; Ghosh:2015opa ; Raju:2015fza ; Gliozzi:2016ysv ; Roumpedakis:2016qcg ; Soderberg:2017oaa ; Behan:2017emf ; Gliozzi:2017hni ; Gliozzi:2017gzh ; Nishioka:2022odm ; Nishioka:2022qmj ; Antunes:2022vtb ; Guo:2023qtt for applications of the multiplet recombination to numerous CFTs, and Safari:2017irw ; Nii:2016lpa ; Hasegawa:2016piv ; Hasegawa:2018yqg ; Skvortsov:2015pea ; Giombi:2016hkj ; Giombi:2017rhm ; Giombi:2016zwa ; Codello:2017qek ; Codello:2018nbe ; Antipin:2019vdg ; Vacca:2019rsh ; Giombi:2020rmc ; Giombi:2020xah ; Dey:2020jlc ; Giombi:2021cnr ; Safari:2021ocb ; Zhou:2022pah ; Bissi:2022bgu ; Herzog:2022jlx ; Giombi:2022vnz ; SoderbergRousu:2023pbe for applications with more emphasis on the equations of motion.
This paper is organized as follows. In section 2, we consider the generalized Yang-Lee edge singularity. We first employ the multiplet recombination to compute the anomalous dimension of $\phi$ and bilinear operators $\mathcal{J}$, then we use the analytic bootstrap to verify that the results are consistent with crossing symmetry. In section 3, we carry out a similar analysis, obtaining the anomalous dimensions in the generalized Potts model, and we check that the results are consistent with crossing symmetry as well. Section 4 focuses on the generalized $\phi^{2n+1}$ theory with $n>1$, where we only consider the single-field case. We determine the anomalous dimensions up to a free parameter, then we verify the consistency between the multiplet-recombination results and crossing symmetry. We summarize the main results in table 1. In appendix A, we work out the action of Laplacians on the scalar three-point function. In appendix B, we verify, in all three cases, that the multiplet-recombination results are consistent with crossing symmetry at subleading twist. We briefly review some representation theory of the symmetric group in appendix C. More details about the tensor structures in the Potts model can be found in appendix D. We compute in appendix E the ratios of OPE coefficients relevant to the multiplet recombination. In appendix F, we derive an OPE coefficient used in the analytic bootstrap. We discuss in appendix G the $N\rightarrow 0$ limit of the generalized Potts model in more detail, especially the logarithmic structures. Appendix H discusses some special limits in the generalized O($N$)-symmetric $\phi^{2n}$ theory.
@@
Table 1: Summary of the main results. The second column lists the anomalous dimensions of $\phi$, and the third column lists the anomalous dimensions of the bilinear operators of the schematic form $\mathcal{J}\sim\phi\,\partial^{\ell}\Box^{m}\phi$.
@@ 2 Generalized Yang-Lee edge singularity
where the upper critical dimension is $d_{\text{u}}=6k$. The parameter $\mu$ has mass dimension 1, so the coupling constant $g$ is dimensionless. We emphasize that $g$ is allowed to be imaginary. Assuming that $\phi$ flips sign under space reflection $\mathcal{P}$ and the unit imaginary number $i$ changes sign under time reversal $\mathcal{T}$, the interaction term is $\mathcal{PT}$ invariant for imaginary $g$. See Bender:2012ea ; Bender:2013qp for more details on the canonical case with $k=1$.
@@ 2.1 Multiplet recombination
$\langle\phi(x_{1})\phi(x_{2})\rangle=\frac{1}{x_{12}^{\smash{2\Delta_{\phi}}}}\,.$ (9)
@@ 2.2 Analytic bootstrap
In this subsection, we verify the consistency between the results above and crossing symmetry. In the lightcone limit of the crossing equation for the four-point function $\langle\phi\phi\phi\phi\rangle$, the anomalous dimensions of the bilinear operators are read off from the terms with enhanced singularity in the crossed channel. [^17] The meaning of the enhanced singularity is as follows. An expression with an enhanced singularity is power-law divergent after repeated action of the conformal Casimir operator (see e.g. Alday:2015eya ). The results are consistent with those derived from the multiplet recombination.
@@ 3 Generalized Potts model
We can extend the $\phi^{3}$ discussion to the case of an $N$-component field $\phi_{a}$, where $a=1,2,\ldots,N$. The action that describes the generalized ($N+1$)-state Potts model is
@@ 3.1 Multiplet recombination
This expression is very similar to (22), but here an $N$-dependent factor appears in the second term in the denominator. [^20] Interestingly, the anomalous dimension of $\phi$ is $k$-independent to order $\epsilon$ at $N=2$: $\gamma_{\phi}=\epsilon/6+O(\epsilon^{3/2})$. For $N=1$, we have $\gamma_{\phi}=O(\epsilon^{3/2})$ due to $T_{111}=0$. See Fang:2022ufx ; Fang:2022tav and references therein for the Fortuin-Kasteleyn representation of the Ising model above the standard upper critical dimension, i.e., $d>4$. At $k=1$, the result agrees with that from the RG analysis Priest:1976zz ; Amit:1976pz . We also obtain $\alpha$ and the OPE coefficient of $\phi$:
@@
For the spin-0 trivial and symmetric traceless operators in the canonical case, the results agree with theumann1984crossover . On the highest trajectory $m=k-1$, the spin-2 current in the trivial representation $(N+1)$ is the stress tensor. As expected, the anomalous dimension of the stress tensor is $O(\epsilon^{3/2})$ according to (69). At $k=1$, the set of leading anomalous dimensions $\{\gamma_{\phi},\gamma_{\mathcal{J}}\}$ contains negative elements when $N>7/3$, which indicates that the canonical Potts model is non-unitary near the upper critical dimension. Moreover, there exists a shadow relation in the standard representation:
$\Delta_{\smash{\phi}}+\Delta_{\mathcal{J}_{\text{standard}}}\big{|}_{\ell=m=0}=d+O(\epsilon^{3/2})\,,$ (71)
which is true for any positive integer $k$. In the $N\rightarrow\infty$ limit, this equation is related to (30) in the generalized Yang-Lee theory. If we take the limit $N\rightarrow 0$, there are degeneracies in the scaling dimensions, leading to logarithmic structures (see appendix G). We also discuss some special limits of the generalized O($N$)-symmetric $\phi^{2n}$ theory in appendix H. In the limit $N\rightarrow-1$, the ratio of OPE coefficients (70) vanishes, but the formal product $R\,\alpha^{2}$ in (69) remains finite because $\alpha^{2}$ diverges as $1/(N+1)^{2}$.
Below, we examine the Chew-Frautschi plots to gain further insights into the theory. According to the procedure in Caron-Huot:2022eqs , the analytic continuation of the trajectories is given by
@@
Figure 1: Chew-Frautschi plot for $\mathcal{J}^{(m=0)}_{\ell,\,\text{trivial}}$ at $N=3$. The dashed lines indicate the $\epsilon\rightarrow 0$ limit of the Regge trajectory. We have set $k=1$ and $\epsilon=10^{-3}$.
Figure 2: Regge trajectories of $\mathcal{J}^{(m=0)}_{\ell,\text{trivial}}$. The $N=\infty$ case corresponds to the decoupled Yang-Lee theory. The plot is made at $k=1$ and $\epsilon=10^{-3}$.
Figure 3: Regge trajectories associated with $\mathcal{J}^{(m=0)}_{\ell,\,\text{trivial}}$. The trajectory in the interacting theory approaches the free theory trajectory as $N\rightarrow 1$. We have used $k=1$ and $\epsilon=10^{-3}$.
@@
$\Delta^{*}_{\pm}-\frac{d}{2}=\pm\left(1+\frac{3R^{\prime}-11N+23}{6(3N-7)}\,\epsilon+O(\epsilon^{3/2})\right)\,,$ (73)
where we have defined $R^{\prime}\equiv R/(N+1)^{2}$. The solution $\Delta^{*}_{-}$ in the standard representation is precisely the scaling dimension of $\phi_{a}$, and $\Delta^{*}_{+}$ is given by $\Delta_{\mathcal{J_{\text{standard}}}}|_{\ell=m=0}$. This is consistent with the shadow relation (71), since $\Delta^{*}_{\pm}$ are related by the shadow transformation $\Delta\rightarrow d-\Delta$.
Figure 4: An $\mathbb{R}^{3}$ projection of the $\mathbb{C}^{2}$ Chew-Frautschi plot for $\mathcal{J}^{(m=0)}_{\ell,\,\text{trivial}}$ at $N=3$. The surfaces are the Riemann surfaces of the functions $\ell(\Delta)$. The imaginary part of $\ell$ is shown by color. Red regions correspond to $\text{Im}[\ell]>0$, and blue regions are associated with $\text{Im}[\ell]<0$. The two turquoise points represent two solutions to $\ell=0$, which are $\Delta_{\pm}$ in (73). The plot is symmetric with respect to the plane $\text{Re}[\Delta-d/2]=0$. We have chosen the parameters $k=1$ and $\epsilon=0.3$.
Figure 5: An $\mathbb{R}^{3}$ projection of the $\mathbb{C}^{2}$ Regge trajectories of $\mathcal{J}^{(m=0)}_{\ell,\,\text{trivial}}$ at $N=3$. The peaks at $\text{Re}[\ell]=-2$ and $\text{Im}[\ell]=0$ correspond to the $1/(\ell+2)$ pole in the anomalous dimension. We indicate the imaginary part of $\Delta-d/2$ by color. The regions with $\text{Im}[\Delta-d/2]>0$ are in red, and the regions with $\text{Im}[\Delta-d/2]<0$ are in blue. Two of the solutions to $\Delta=d/2$ are $\ell_{\pm}$ in (74). They form a complex conjugate pair, and are indicated by the two turquoise points. We have used $k=1$ and $\epsilon=0.3$.
Figure 6: A 3D Chew-Frautschi plot for $\mathcal{J}^{(m=0)}_{\ell,\,\text{trivial}}$ at $N=6/5$. Color coding indicates the imaginary part of $\ell$: red for $\text{Im}[\ell]>0$ and blue for $\text{Im}[\ell]<0$. The turquoise points are the solutions to $\ell=0$, given by $\Delta_{\pm}$ in (73). The plot is symmetric relative to the plane $\text{Re}[\Delta-d/2]=0$. The parameters are $k=1$ and $\epsilon=0.3$.
@@
Figure 7: An $\mathbb{R}^{3}$ projection of the $\mathbb{C}^{2}$ Chew-Frautschi plot for $\mathcal{J}^{(m=0)}_{\ell,\,\text{trivial}}$ at $N=6/5$. The peaks at $\text{Re}[\ell]=-2$, $\text{Im}[\ell]=0$ is due to the $1/(\ell+2)$ pole in the anomalous dimension. The imaginary part of $\ell$ is shown by color: red regions implies $\text{Im}[\ell]>0$, and blue regions indicate $\text{Im}[\ell]<0$. The turquoise points mark the solutions $\ell_{\pm}^{*}$ in (74). The plot is made at $k=1$ and $\epsilon=0.3$.
@@
The fact that the leading real intercept is negative should have interesting physical implications. For instance, in the Regge limit, the connected part of the four-point correlator decays exponentially fast at large boost if the Regge intercept is below unity Costa:2012cb ; Caron-Huot:2017vep ; Liu:2020tpf ; Caron-Huot:2020ouj ; Caron-Huot:2022eqs . [^21] The anomalous dimension (69) is regular at $\ell=0$, which corresponds to a local spin-0 operator. The analytic continuation above leads to both the local operator and its shadow, but the analytic continuation in conformal spin only gives the local operator. This is similar to the case of the $\phi^{2n}$ theory with $n>2$ Guo:2023qtt .
@@
Now, we briefly comment on the $k>1$ Chew-Frautschi plots. One difference is the presence of more curves, which arise from the deformation of more horizontal lines in the Gaussian limit. Another difference is that we can consider trajectories with higher $m$. For each $m$, there are more solutions to $\Delta=d/2$, so the Regge trajectory can have more intersections with the vertical axis. Let us first discuss the Gaussian limit, where the scaling dimensions do not depend on the $S_{N+1}$ representations. For a positive integer $k$, the horizontal lines are given by $\frac{(1-3k-\ell-m)_{2k}}{\ell+k+2m}=0$, implying that there are $12k-$ 1 such lines. The diagonal lines are translations of those in the $k=1$ case. In the interacting theory, the trajectories are deformed into disconnected pieces. For example, the $\mathcal{J}^{(m=0)}_{\ell,\,\text{trivial}}$ trajectory in the $k=2$ theory and its Gaussian limit are shown in figure 8. As in the $k=1$ case, the Chew-Frautschi plots in the interacting theory depend on $N$. For both $m=0$ and $m=1$, there are three scenarios: $N>65/37$, $371<N<65/$ 37 and $N<1$. We present the Chew-Frautschi plots with different values of $N$ in figures 9, 10, 11 and 12. The Chew-Frautschi plots with $N<1$ are similar to case of $N>65/37$.
Figure 8: Chew-Frautschi plot for $\mathcal{J}^{(m=0)}_{\ell,\,\text{trivial}}$ in the $k=2$ theory. The interacting case with $N=2$ is shown in red. In the Gaussian limit, there are two dashed diagonal lines given by $(\ell+2)^{2}=(\Delta-6)^{2}$, and there exist three dashed horizontal lines at $\ell=-3,-4,-5$.
Figure 9: In the $k=2$ theory, the Regge trajectories of $\mathcal{J}^{(m=0)}_{\ell,\,\text{trivial}}$ are shown. These trajectories correspond to the case of $N>65/37$. We have set $\epsilon=10^{-3}$.
Figure 10: Chew-Frautschi plot of the $k=2$ interacting theory. The Regge trajectories of $\mathcal{J}^{(m=0)}_{\ell,\,\text{trivial}}$ with $371<N<65/$ 37 are shown. We have used $\epsilon=10^{-3}$.
Figure 11: Regge trajectories of $\mathcal{J}^{(m=1)}_{\ell,\,\text{trivial}}$ in the $k=2$ interacting theory. We present the scenario where $N>65/37$. Again, $\epsilon=10^{-3}$ is used.
Figure 12: In the $k=2$ interacting theory, we show the Regge trajectories of $\mathcal{J}^{(m=1)}_{\ell,\,\text{trivial}}$ with $371<N<65/$ 37, where $\epsilon$ is set to $10^{-3}$.